OpenStudy (anonymous):

Can someone check my answers? 11 – 4x = 5 I got 1.2925 7e^2x - 5 = 27 I got 0.760 5x - 1 = 3 I got 2000 14^x = 5 I got 0.6099

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hey

OpenStudy (misty1212):

put the calculator away! \[11-4x=5\\ -4x=5-11\\ -4x=-6\\ x=\frac{-6}{-4}\\ x=\frac{3}{2}\]

OpenStudy (anonymous):

it's 4^x

OpenStudy (misty1212):

i now am guessing that there are exponents there use the equation tool, it is much clearer

OpenStudy (misty1212):

\[11-4^x=5\]?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

\[4^x=6\]

OpenStudy (misty1212):

\[x=\frac{\ln(6)}{\ln(4)}\]